### v2

# A battle of designs: triangular vs. L-shaped detectors and parity violation in the gravitational-wave background

Abstract
We investigate the prospects for detecting a parity-violating gravitational-wave background (GWB) with third-generation ground-based detector networks. We focus on a network consisting of one Einstein Telescope (ET) and two Cosmic Explorer (CE) detectors. In our analysis we vary the ET design, detector orientations, and arm lengths to assess the impact of geometry and scale on detection capabilities. We find that, among the configurations studied, networks with a $15$ km triangular ET design provide the strongest parity-violation constraining power, followed by networks with an L-shaped ET design, while networks with $10$ km triangular ET configurations are the least sensitive. Under current bounds, ET alone cannot confidently detect parity violation.
preprint: KCL-PH-TH-2025-44
I Introduction
The superposition of independent gravitational waves (GW) of cosmological and astrophysical origin forms a gravitational-wave background (GWB). The detection of such a background can shed light on early-universe mechanisms, such as inflation Grishchuk (1974); Starobinsky (1979); Badger et al. (2024), first-order phase transitions Kamionkowski et al. (1994a); Apreda et al. (2002); Grojean and Servant (2007); Romero et al. (2021); Badger and others (2023), and cosmic strings Auclair and others (2020); Damour and Vilenkin (2001); Abbott and others (2021), and aid in characterizing astrophysical compact binary coalescences (CBCs) sources and populations Abbott and others (2023). Searches for an unpolarized GWB using data from the ground-based interferometers LIGO Aasi and others (2015) and Virgo Acernese and others (2015) have placed upper limits on its energy density Abac and others (2025b).
A multitude of mechanisms in the early universe can generate parity violation (PV) Alexander et al. (2006) that results in the production of asymmetric amounts of right- and left-handed circularly polarized GWs. Parity-violating sources and their effects on the GWB have been studied in the literature, including those arising from the Chern-Simons modification to gravity Satoh et al. (2008); Bartolo et al. (2021); Takahashi and Soda (2009) and from axion inflation Barnaby and Peloso (2011); Dimastrogiovanni and Peloso (2013), turbulence in the primordial plasma Kamionkowski et al. (1994b); Witten (1984); Hogan (1986), or by primordial magnetic fields coupled to the cosmological plasma Brandenburg et al. (1996); Christensson et al. (2001); Kahniashvili et al. (2010); Brandenburg et al. (2019, 2021) .
Probing for polarization in the GWB can place constraints on parity-violating theories Crowder et al. (2013). Parity-violating effects in the GWB have been studied in the context of current and future detector sensitivities Seto (2007); Ricciardone (2017); Orlando et al. (2021); Bartolo et al. (2018); Domcke et al. (2020); Xu et al. (2019); Martinovic et al. (2021a); Badger and Sakellariadou (2021); Cruz et al. (2024), and have been constrained through analyses of LIGO–Virgo data Crowder et al. (2013); Martinovic et al. (2021a); Abac and others (2025a).
While PV detection appears unlikely in presently existing ground detectors, third-generation detectors show promise in detecting, or at least imposing tighter constraints in circular polarization. The choice of the detector network configuration will play a crucial role in the circular polarization detection prospects. In this study, we present a dual theoretical (based on signal-to-noise ratio) and data-driven (Bayesian) analysis for assessing the sensitivity of third-generation ground-based detector networks to parity-violating signatures in a general power-law GWB model.
II Formalism
The right- and left-handed modes, defined through the circularly polarized bases $e^{\rm{R}}=(e^{+}+ie^{\times})/\sqrt{2}$ and $e^{\rm{L}}=(e^{+}-ie^{\times})/\sqrt{2}$ (with + and $\times$ the plus and cross polarizations, respectively) read $h_{\rm R}=(h_{+}-ih_{\times})/\sqrt{2}$ and $h_{\rm L}=(h_{+}+ih_{\times})/\sqrt{2}$.
Right- and left-handed correlators are
$\begin{pmatrix}\langle h_{\rm R}(f,\hat{\Omega})h_{\rm R}^{*}(f^{\prime},\hat{\Omega}^{\prime})\rangle\\ \langle h_{\rm L}(f,\hat{\Omega})h_{\rm L}^{*}(f^{\prime},\hat{\Omega}^{\prime})\rangle\end{pmatrix}$ (1)
[…]
where $\langle\cdot\rangle$ is the ensemble average and $I,V$ are the Stokes parameters; $V$ characterizes the asymmetry between right- and left-handed polarized waves, and $I(\geq|V|)$ is the gravitational wave’s total amplitude. The polarization degree $\Pi(f)=V(f)/I(f)\in[-1,1]$, with $\Pi=-1$ and $\Pi=+1$ defining fully left- and right-handed polarization, respectively, while $\Pi=0$ stands for an unpolarized isotropic GWB. For the latter ($V=0$), Eq. (1) reduces to the standard correlator for an unpolarized isotropic background. The signal-to-noise ratios (SNRs) $\rho_{I}$ and $\rho_{V}$ for the $I$- and $V$-modes, respectively, of a circularly polarized GWB, are calculated for a detector network over an observing time $T$ Seto and Taruya (2008). The power-law integrated (PI) curve, $\Omega_{{\rm{PI}},I}$, characterizes a detector network’s $I$-mode sensitivity Thrane and Romano (2013). We extend the procedure to the $V$-mode to obtain the PI sensitivity curve, $\Omega_{{\rm{PI}},V}$, for the $V$-mode.
To search for parity-violating GWBs, we adapt the formalism described in  Seto and Taruya (2008, 2007) to third-generation ground-based detector networks.
III Parity Violating Models
We explore the parity-violation detection prospects of the modified GWB spectrum from formalism Seto and Taruya (2007); Crowder et al. (2013)
$\Omega_{\rm GW}^{\prime}(f)=\Omega_{\rm GW}(f)\Bigg[1+\Pi(f)\frac{\gamma_{V}^{d_{1}d_{2}}(f)}{\gamma_{I}^{d_{1}d_{2}}(f)}\Bigg]~,$ (2)
as it would appear in the detector pairs $d_{1}$, $d_{2}$ where $\gamma_{I/V}^{d_{1}d_{2}}$ are the normalized overlap reduction functions for the $I$- and $V$-modes calculated via analytical and numerical methods Romano and Cornish (2017); Caporali and Ricciardone (2025) defined as the usual
$\gamma_{I}^{d_{1}d_{2}}(f)=\frac{5}{8\pi}\int d\hat{\Omega}(F_{d_{1}}^{+}F_{d_{2}}^{+*}+F_{d_{1}}^{\times}F_{d_{2}}^{\times*})e^{2\pi if\hat{\Omega}\Delta\overrightarrow{x}}~,$ (3)
$\gamma_{V}^{d_{1}d_{2}}(f)=-\frac{5i}{8\pi}\int d\hat{\Omega}(F_{d_{1}}^{+}F_{d_{2}}^{\times*}-F_{d_{1}}^{\times}F_{d_{2}}^{+*})e^{2\pi if\hat{\Omega}\Delta\overrightarrow{x}}~,$ (4)
for beam pattern function $F^{A}_{d}$ for the usual $A=+,\times$ polarizations.
We assume a general power-law GWB spectrum $\Omega_{\rm{GW}}=\Omega_{\alpha}(f/f_{\rm{ref}})^{\alpha}$ parametrized by amplitude $\Omega_{\alpha}$ and spectral index $\alpha$ Martinovic et al. (2021a); we set the reference frequency $f_{\rm ref}=25~\rm{Hz}$. Theoretically motivated polarized GWBs can be modeled by specific choices of $\alpha$, with the polarization typically assumed constant for simplicity. For instance, GWBs sourced by non-Abelian axion inflation  Badger et al. (2024); Maleknejad (2016), string-inspired cosmology models Satoh et al. (2008), and some turbulence models Kahniashvili et al. (2005) can be well-modeled by a power-law GWB with constant polarization. In what follows, we adopt this assumption, $\Pi(f)=\rm{const}$, noting that the framework can be generalized to arbitrary polarized GWBs, for example, a frequency-dependent power-law $\Pi(f)$ model Martinovic et al. (2021a).
IV Future ground-based detector configurations
We consider seven representative networks, each combining one Einstein Telescope (ET) with two Cosmic Explorer (CE) detectors, using the geometric information from  Ebersold et al. (2024). We include triangular ET designs with $10$ and $15$ km arms ($\Delta_{10}$, $\Delta_{15}$), as well as ET designs consisting of two L-shaped detectors with $15$ km arms (S, R), following the standard ET configurations introduced in  Branchesi and others (2023) and extended to ET+CE networks in  Ebersold et al. (2024). For CE, we consider L-shaped detectors with $20$ km (P) and $40$ km (Y) arms. The triangular configurations consist of three detectors, while an L-shaped configuration consists of one. Thus a $\Delta$ + (PY) network contains five detectors ($10$ baselines), while an (SR) + (PY) network contains four ($6$ baselines). These configurations explore how detector orientations and arm length influence network performance, with some optimized for GWB sensitivity and others for CBC source localization, clarifying whether differences between triangular and L-shaped designs arise from geometry or arm length. Further technical details are in Ebersold et al. (2024), for which we use the amplitude spectral density curves found in Danilishin and Zhang (2023) and Fulda et al. (2023), and the full set of configuration names and descriptions is in Table 1.
Network name Description
$\Delta_{10}$(PY)4 $10$ km $\Delta$ + 2CE, GWB-optimized
$\Delta_{10}$(PY)5 $10$ km $\Delta$ + 2CE, Localization-optimized
$\Delta_{15}$(PY)4 $15$ km $\Delta$ + 2CE, GWB-optimized
$\Delta_{15}$(PY)5 $15$ km $\Delta$ + 2CE, Localization-optimized
(SR)1(PY)1 $15$ km 2L + 2CE, GWB-optimized
(SR)2(PY)2 $15$ km 2L + 2CE, Localization-optimized
(SR)3(PY)3 $15$ km 2L + 2CE, Hybrid
Table 1: The different third-generation detector network configurations that are considered in this study, which follow the definitions and optimization procedures of Ebersold et al. (2024). “GWB-optimized” denotes configurations favoring GWB sensitivity and are determined using a PI-curve method, while “Localization-optimized” configurations favor source localization and are computed using the network-alignment factor. “Hybrid” configurations provide a balanced design providing a trade-off between the two.
In Fig. 1, we present the $I$- and $V$-mode PI curves for the various network configurations, extending previous ET design studies Branchesi and others (2023) to networks combining ET with CE and to $V$-mode sensitivity. The $\Omega_{{\rm{PI}},I}$ curves show that GWB-optimized networks yield the highest $I$-mode sensitivity, which is expected by construction. In particular, (SR)1(PY)1 dominates below $20$ Hz, while at higher frequencies all triangular configurations outperform the 2L designs.
The $\Omega_{{\rm{PI}},V}$ curves exhibit reduced sensitivity compared to their $I$-mode counterparts, with sensitivity decreasing as $|\Pi|$ approaches 0. Their ranking is relatively stable across frequencies: $\Delta_{15}$(PY)4 is consistently the most sensitive configuration, while $\Delta_{10}$(PY)5 performs the worst. Overall, networks incorporating a $15$ km triangular ET design achieve the highest sensitivity to $V$-modes, followed by 2L ET networks, whereas $10$ km triangular ET configurations are the least sensitive. The superior performance of the $15$ km triangular configuration relative to the $15$ km 2L case arises from its larger number of independent baselines. In the $10$ km triangular configuration, however, higher instrumental noise dominates, leading to reduced sensitivity. Sensitivity to $V$-modes is therefore controlled by a combination of arm length, the number of independent baselines, and network geometry. The parity-violation detection prospects of ET alone are presented in Appendix  A. We find that current constraints Abac and others (2025a) have already excluded the parameter space in which ET alone would be capable of detecting circular polarization. Therefore, the inclusion of CE detectors is a necessity to make robust parity-violation inferences.
Figure 1: PI sensitivity curves for $I$-mode (solid lines) and $V$-mode (dotted and dashed lines), shown for the seven different detector third-generation network configurations for an observing time of one year and $\rho=4$. The $V$-mode curves correspond to $|\Pi|=1$ (dotted) and $|\Pi|=0.1$ (dashed).
V Future detection prospects
We present theoretical (based on signal-to-noise ratio) and data-driven (obtained with a Bayesian analysis) prospects for detecting PV by applying the previously described formalisms to third-generation detector networks. Our analysis covers general power-law GWBs with $12.5\leq\log_{10}\Omega_{\alpha}\leq-10$, spectral indices $\alpha=-3,\dots,3$, and polarization degrees $1\leq\Pi\leq 1$. We note that the upper bound on the $\log_{10}\Omega_{\alpha}$ range is determined by current detector constraints Abac and others (2025b), while the lower bound is to allow studied GWBs to be potentially detected in the considered networks, as seen in Fig. 1.
V.1 Signal-to-noise ratio
The sensitivities of the third-generation detector networks are quantified by the $I$- and $V$-mode SNRs. Fig. 2 shows the results for a flat PL ($\alpha=0$) spectrum across the seven network configurations. As expected, $\rho_{I}$ is independent of $\Pi$ and scales only with the GWB amplitude, while $\rho_{V}$ grows with $|\Pi|$, enhancing detectability for more strongly polarized backgrounds. We repeat this study for other power-law indices and rank their detection prospects in Tables 3 and 3.
Figure 2: SNRs for $I$-mode (solid lines) and $V$-mode (dashed lines) for the seven third-generation detector network configurations, assuming a flat ($\alpha=0$) power-law GWB.
Order $\alpha$ 3,2-3,-2 $1,0,\tfrac{1}{3},\tfrac{1}{2},\tfrac{2}{3},1$ 22 33
1st (SR)1(PY)1 (SR)1(PY)1 $\Delta_{15}$(PY)4
$\Delta_{15}$(PY)4
$\Delta_{15}$(PY)5
2nd $\Delta_{15}$(PY)4 $\Delta_{15}$(PY)4 (SR)1(PY)1
$\Delta_{10}$(PY)4
$\Delta_{10}$(PY)5
3rd $\Delta_{15}$(PY)5 (SR)3(PY)3 $\Delta_{15}$(PY)5 (SR)1(PY)1
4th (SR)3(PY)3 $\Delta_{10}$(PY)4 $\Delta_{10}$(PY)4 (SR)3(PY)3
5th $\Delta_{10}$(PY)4 $\Delta_{15}$(PY)5 (SR)3(PY)3 (SR)2(PY)2
6th (SR)2(PY)2 $\Delta_{10}$(PY)5 $\Delta_{10}$(PY)5
7th $\Delta_{10}$(PY)5 (SR)2(PY)2 (SR)2(PY)2
Table 2: Ranking from best (1st) to worst (7th) of detector networks by II-mode SNRs for different PL indices $\alpha$, grouping together indices that yield identical prospects.
Order $\alpha$ 3-3 to 11 $2,3$
1st $\Delta_{15}$(PY)4 $\Delta_{15}$(PY)4
2nd $\Delta_{15}$(PY)5 $\Delta_{15}$(PY)5
3rd (SR)3(PY)3 (SR)1(PY)1
4th (SR)1(PY)1 (SR)3(PY)3
5th (SR)2(PY)2 (SR)2(PY)2
6th $\Delta_{10}$(PY)4 $\Delta_{10}$(PY)4
7th $\Delta_{10}$(PY)5 $\Delta_{10}$(PY)5
Table 3: Ranking from best (1st) to worst (7th) of detector networks by $V$-mode SNRs for different PL indices $\alpha$.
We find that the 2L network (SR)1(PY)1 provides the largest $\rho_{I}$ for $3\leq\alpha\leq 1$, while the $15$ km triangular network $\Delta_{15}$(PY)4 overtakes it at $\alpha=2$, highlighting the advantage of a triangular design for GWBs with large positive spectral indices. For the $V$-mode, $\Delta_{15}$(PY)4 is the most sensitive configuration across all $\alpha$ values, with the 2L networks consistently outperforming the $10$ km triangular configurations. This hierarchy indicates that optimal parity-violation detection is achieved with a $15$ km triangular ET combined with two CEs. This behavior is consistent with the PI curves in Fig. 1, whose frequency-dependent shapes qualitatively explain the SNR rankings across different spectral indices. We highlight that fixing different reference frequencies yields the same ordering of SNR prospects, emphasizing the generality of our results. Appendix A presents ET-only configurations and shows that all triangular ET designs suffer a dramatic loss of PV sensitivity compared to 2L designs.
V.2 Bayesian analysis
We now complement the SNR analysis with a data-driven Bayesian analysis on simulated signals, assuming the design sensitivity and noise characteristics of each third-generation network configuration. Parameter estimation is carried out with the pygwb Renzini and others (2024) and Bilby  Ashton and others (2019) packages using the Dynesty sampler. This allows us to assess a network’s capability to detect a polarized signal and to distinguish it from an unpolarized one. To this end, we fix a power-law index $\alpha$ and generate 1600 simulated injections uniformly over the $\log_{10}\Omega_{\alpha}$ and $\Pi$ range in each network configuration, then searching over the priors of Table 4 based off of theoretical bounds and previously explored search priors Martinovic et al. (2021b); Abac and others (2025a). The procedure is then repeated with $\Pi=0$ to model unpolarized backgrounds, using the same amplitude and spectral index priors. To reduce statistical noise, we generate $15$ independent injection sets with random detector noise and combine results with an optimized weighted average Romano and Cornish (2017). This enables a direct comparison between polarized and unpolarized searches.
Parameters $\theta$ Priors
$\log_{10}\Omega_{\rm\alpha}$ Uniform(18,7-18,-7)
$\alpha$ Gaussian($0,3.5$)
$\Pi$ Uniform(1,1-1,1)
Table 4: Priors used in the Bayesian analysis. The $\log_{10}\Omega_{\rm\alpha}$ and $\alpha$ priors follow previous searches, while taking into account the sensitivity of third-generation detectors Abbott et al. (2021); Martinovic et al. (2021b). The prior on $\Pi$ is the full, physically allowed range of values Martinovic et al. (2021a).
V.2.1 Comparison to signal-to-noise
We compare the recovered data-analysis Bayes factors to the $I$- and $V$-mode SNRs for all injections. Fig. 3 shows this comparison for the (SR)3(PY)3 and $\Delta_{10}$(PY)5 configurations for a flat GWB spectrum. More concretely, it shows the SNRs alongside the log Bayes factors for a polarized model compared to detector noise, $\ln\mathcal{B}_{\rm Noise}^{\Pi\neq 0}$, and to an unpolarized model, $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}$. We adopt strong detection thresholds of $\ln\mathcal{B}_{\rm Noise}^{\Pi\neq 0}$ and $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}$ equal to 8. These correspond to $\rho_{I}$ and $\rho_{V}$ equal to 4, since $\rho^{2}\simeq 2\ln\mathcal{B}$ in the strong signal regime Romano and Cornish (2017). Our findings confirm this relation.
Figure 3: SNR and log Bayes detection prospects for II- and $V$-modes in a flat GWB in the $\Delta_{10}$(PY)5 and (SR)3(PY)3 configurations. The green and purple filled regions correspond to $95\%$ exclusion regions of $\Pi=0$. The gray-shaded region indicates where $\rho_{V}^{\Delta_{10}\text{(PY)}_{5}}\leq\rho_{V}^{95\%,~\rm{crit}}\leq\rho_{V}^{\text{(SR)}_{3}\text{(PY)}_{3}}$, where $\rho_{V}^{95\%,~\rm{crit}}=2.16$ is the approximate $V$-mode SNR where $\Pi=0$ is excluded at 95%. The red star marks the case study discussed in Sec. V.2.2.
In conclusion, the SNR predictions reproduce the trends seen in the data-driven analysis across both networks, demonstrating that simple theoretical estimates based on SNRs provide an excellent guide to the detectability of polarized GWBs. We note that the present analysis assumes uncorrelated detector noise. For the ET triangular configuration, correlated Newtonian noise between the co-located detectors is expected to further degrade the sensitivity to GWBs Janssens and others (2024). Since such correlated noise is absent for spatially separated L-shaped detectors, including this effect would reduce the performance of the triangular network and strengthen the preference for 2L configurations found here.
V.2.2 Parameter estimation
We perform a Bayesian analysis to assess parameter recovery of simulated parity-violating GWBs for the (SR)3(PY)3 and $\Delta_{10}$(PY)5 configurations. From the marginalized posteriors, we compute the 95% Upper Limits (UL) and Lower Limits (LL) on $\Pi$, allowing us to make $\Pi=0$ exclusion analyses. This is shown in Fig. 3: purple and green regions denote where $\Pi=0$ is ruled out with at least $95\%$ confidence by the 2L and triangular networks, respectively. We define a threshold region as the parameter space where the 2L network excludes $\Pi=0$ at $95\%$ confidence, but the triangle network does not. An average $V$-mode SNR corresponding to the borders of this threshold region is $\rho_{V}^{\Delta_{10}\text{(PY)}_{5}}\leq 2.16\leq\rho_{V}^{\text{(SR)}_{3}\text{(PY)}_{3}}$, which is indicated by the gray-shaded region in Fig.  3.
Figure 4: Posterior distributions from Bayesian inference for parity-violating power-law GWBs, combined over $15$ simulated runs for $f_{\rm ref}=25$ Hz. Shown are the (SR)3(PY)3 (purple) and $\Delta_{10}$(PY)5 (green) networks, with dark and light shading indicating $1\sigma$ and $2\sigma$ credible regions. Here: $\Pi=-1$ and $\log_{10}\Omega_{\alpha}=-11.54$ shown in blue.
Next, we focus on a flat parity-violating GWB case study with $\log_{10}\Omega_{\alpha}=-11.54$, $\Pi=-1$, indicated by the red star in Fig. 3. These parameters were chosen so that the signal lies in the threshold regime. The combined posterior results are shown in Fig. 4, where model parameters are recovered within $1\sigma$ for both networks. We derive the $95\%$ UL on $\Pi$ equal to 0.238-0.238 and $0.233$ for the 2L and triangular network, respectively. We also compute log Bayes factors for the parity-violating and the unpolarized GWB searches. For the (SR)3(PY)3 network, we obtain $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}=2.914\pm 0.043$, while for $\Delta_{10}$(PY)5 we find $\ln\mathcal{B}_{\Pi=0}^{\Pi\neq 0}=0.946\pm 0.037$, highlighting the greater capability of (SR)3(PY)3 to distinguish a polarized from unpolarized GWBs than $\Delta_{10}$(PY)5. Although these log Bayes factors do not indicate a strong preference for a PV model, this injection demonstrates that (SR)3(PY)3 can rule out $\Pi=0$ at the $95\%$ confidence level, while $\Delta_{10}$(PY)5 cannot.
In conclusion, these findings highlight that network design strongly impacts parameter constraints, particularly for circular polarization $\Pi$, with (SR)3(PY)3 providing tighter constraints on parity-violating GWBs than $\Delta_{10}$(PY)5.
VI Conclusions
In this study, the detection prospects of a general parity-violating GWB with third-generation detector networks were investigated using a signal-to-noise ratio and a Bayesian analysis approach. By extending the standard GWB formalism to include the $V$-mode, we introduced PI curves and SNR prospects that accurately reproduce data-based results, providing an efficient proxy for assessing circular polarization detection prospects.
We considered an ET detector, in both triangular and 2L designs, paired with two CE detectors across a range of orientations, and ET arm lengths. A broad set of power-law GWB spectra, motivated by cosmological parity-violating mechanisms, was analyzed.
Across all scenarios, $15$ km triangular ET configurations are most sensitive to parity violation, followed by 2L configurations, while $10$ km triangular ET configurations are the least sensitive. Parameter estimation studies further demonstrate that network design strongly affects the ability to constrain the GWB parameters. We find that adopting a more sensitive network configuration can tighten the searched parameter space and can place polarization bounds that exclude $\Pi=0$ at 95% compared to less sensitive detector designs. In Appendix A, we show that parity violation detection in ET alone has been ruled out; circular polarization in a third-generation network requires CE detectors.
These findings underscore the importance of a global network design choices in order to optimize the discovery potential for fundamental physics in the GWB. These results highlight how next-generation detectors could open a direct observational window onto fundamental parity-violating physics in the early Universe.
Acknowledgements.
We are grateful to Jacopo Uggeri for extensive discussions. We also thank Michael Ebersold for valuable insights on third-generation network configurations, Alberto Mariotti and Miguel Vanvlasselaer for helpful discussions, and Gianmassimo Tasinato and Angelo Ricciardone for useful comments. This document has an ET document number ET-0483A-25. The work of MS is partially supported by the Science and Technology Facilities Council (STFC grant ST/X000753/1). HD is supported by the Strategic Research Program High-Energy Physics of the Research Council of the Vrije Universiteit Brussel, the iBOF “Unlocking the Dark Universe with Gravitational Wave Observations: from Quantum Optics to Quantum Gravity” of the Vlaamse Interuniversitaire Raad.
Appendix A Discussion on Einstein Telescope alone
We briefly assess the detection prospects of a parity violating GWB using ET alone, building on previous ET-only detector design and sensitivity studies Branchesi and others (2023). For this purpose, we consider baselines that differ from those adopted in the main analysis. Specifically, we introduce the notation $\Delta^{\prime}_{10}$ and $\Delta^{\prime}_{15}$ to denote closed, fully symmetric triangular ET configurations with $10$ km and $15$ km arms, respectively. These correspond to the idealized case of a perfectly equilateral triangle formed by three co-located detectors, with each vertex separated by 60°. For comparison, we also include the configurations $\Delta_{10}$ and $\Delta_{15}$ already used in the main text and defined in Ebersold et al. (2024). These differ in that the three detectors do not form an exactly closed triangle. As a result, the $\Delta_{10}$ and $\Delta_{15}$ configurations are slightly asymmetric. This asymmetry yields enhanced sensitivity to $V$-modes, since the chiral ORF is slightly larger than that of a perfectly closed triangle.
In addition to triangular configurations, we also examine two 2L configurations, denoted by (SR)0 and (SR)45. The (SR)0 setup corresponds to the optimal 2L configuration for GWB detection where the two detectors are aligned (relative orientation of 0°). By contrast, the (SR)45 represents the most pessimistic case for GWB detection, where the detectors are rotated by 4545° with respect to one another, consistent with previous studies of 2L configurations and relative orientations Branchesi and others (2023)). It is important to emphasize that (SR)0 and (SR)45 differ from (SR)1 and (SR)2 considered in the main analysis, since the latter were defined relative to the optimal configuration of a global network including two CEs, whereas here the reference is ET alone. In Table 5, the latitude and longitude of each network center, the X- and Y-arm endpoints and azimuths are given.
Name Origin (lat, lon) X-arm end (lat, lon) Y-arm end (lat, lon) Azimuth [°] X-/Y-arm
E1’10 43.6300, 10.5000 43.7149, 10.5413 43.6983, 10.4193 19.43 /319.43
E2’10 43.7149, 10.5413 43.6983, 10.4193 43.6300, 10.5000 259.46 /199.46
E3’10 43.6983, 10.4193 43.6300, 10.5000 43.7149, 10.5413 139.37 /79.37
E1’15 43.6300, 10.5000 43.7573, 10.5620 43.7325, 10.3789 19.43 /319.43
E2’15 43.7573, 10.5620 43.7325, 10.3789 43.6300, 10.5000 259.47 /199.47
E3’15 43.7325, 10.3789 43.6300, 10.5000 43.75730, 10.5620 139.35/79.35
S0 40.5200, 9.4170 40.3851, 9.4093 40.5258, 9.2402 182.51/272.51
S45 40.5200, 9.4170 40.4203, 9.5364 40.4287, 9.2867 137.51 /227.51
R0=R45 50.7200, 5.9210 50.8548, 5.9210 50.7198, 6.1334 0.00/90.00
Table 5: ET coordinates and azimuths (clockwise from North).
Fig. 5 shows the PI curves for both the $I$- and $V$-modes across the different ET-only configurations. For the $I$-mode sensitivities, the trends are consistent with previous ET-only sensitivity studies Branchesi and others (2023), while here we additionally present the corresponding $V$-mode sensitivities relevant for parity-violating GWBs. For comparison, we also include a parity-violating PL GWB with spectral index $\alpha=0$, to facilitate direct comparison with the corresponding SNR prospects shown in Fig. 6. The perfectly closed and symmetric triangular configurations $\Delta^{\prime}$ exhibit noticeably lower $V$-mode sensitivity than the slightly asymmetric $\Delta$ setups used in the main text, illustrating that even small asymmetries in ET’s geometry can enhance parity-violation detection. Nevertheless, all triangular configurations, whether symmetric or slightly asymmetric, are significantly less sensitive to the $V$-modes than the 2L designs.
Figure 5: PI curves sensitivity curves for $I$-mode (solid lines) and $V$-mode (dotted and dashed lines), shown for the 6 different ET configurations, for an observing time of one year and $\rho=4$. The $V$-mode curves correspond to $|\Pi|=1$ (dotted) and $|\Pi|=0.1$ (dashed). We also show PL GWBs with spectral index of $\alpha=0$.
Figure 6: SNRs for $I$-mode (solid lines) and $V$-mode (dashed lines), shown for the five different ET configurations for a PL GWB with $\alpha=0$.
We have again performed SNR studies for multiple spectral indices $\alpha$. For the $I$-mode SNRs, the ranking from most to least sensitive is: (SR)0, (SR)45, $\Delta_{15}=\Delta_{15}^{\prime}$ and $\Delta_{10}=\Delta_{10}^{\prime}$. For the $V$-mode SNRs, the resulting network rankings are summarized in Tables 6, confirming that 2L configurations consistently outperform triangular ET for $V$-mode detection. We note that correlated Newtonian noise between the co-located detectors of the triangular configuration Janssens and others (2024) is not included in the present study. Accounting for this contribution would further reduce the sensitivity prospects of the triangular ET designs. Moreover, current bounds on the parity-violated GWB Abac and others (2025a) have ruled out the parity-violating parameter space that is detectable by both 2L and the $\Delta$ configurations, as seen in Fig.6. Consequently, if ET were constructed without CE, none of the configurations would be able to detect circularly polarized GWBs at a confident SNR level.
Order $\alpha$ 3,2,1-3,-2,-1 $0,\tfrac{1}{2},1,\tfrac{2}{3},2,3$
1st (SR)45 (SR)0
2nd (SR)0 (SR)45
3rd $\Delta_{15}$ $\Delta_{15}$
4th $\Delta_{10}$ $\Delta_{10}$
5th $\Delta_{15}^{\prime}$ $\Delta_{15}^{\prime}$
6th $\Delta_{10}^{\prime}$ $\Delta_{10}^{\prime}$
Table 6: Ranking of ET configurations from most (1st) to least (6th) sensitive for different spectral indices $\alpha$ for the $V$-mode SNRs.